# Even more efficient magic state distillation by zero-level distillation

### v1

@@ Abstract
Magic state distillation (MSD) is an essential element for universal fault-tolerant quantum computing, which distills a high fidelity magic state from noisy magic states using ideal (error-corrected) Clifford operations. For ideal Clifford operations, it needs to be performed on the logical qubits and hence takes a large spatiotemporal overhead, which is one of the major bottlenecks for the realization of fault-tolerant quantum computers (FTQC). Here we propose zero-level distillation, which prepares a high fidelity logical magic state using physical qubits on a square lattice using nearest-neighbor two-qubit gates without using multiple logical qubits. The key idea behind is using the Steane code to distill a logical magic state by using noisy Clifford gates with error detection. Then the Steane code state is teleported or converted to the surface codes. By carefully designing such circuits fault-tolerantly, the error rate of the logical magic state scales $\sim 100\times p^{2}$ in terms of the physical error rate $p$. For example, with a physical error rate of $p=10^{-4}$ ($10^{-3}$), the logical error rate is reduced to $p_{L}=10^{-6}$ ($10^{-4}$), resulting in an improvement of two (one) orders of magnitude. This contributes to reducing both space and time overhead for early FTQC as well as full-fledged FTQC combined with conventional multi-level distillation protocols.
@@ I Introduction
Quantum computers are expected to provide advantages in solving problems that are intractable for classical computers such as prime factorization [1], linear system solver [2], and quantum chemistry [3]. Significant experimental efforts have been dedicated to the realization of quantum computers based on various physical systems. Among these, the superconducting system is one of the most promising candidates, and systems with 50-100 qubits have already been experimentally demonstrated [4, 5]. These quantum computers are called noisy intermediate-scale quantum computers (NISQ) [6], whose noise level is still high and the number of qubits is still limited, are currently unable to run sophisticated quantum algorithms with theoretically proven quantum speedup. Although NISQ-aware algorithms are being developed [7], an ultimate solution to these problems is to protect quantum information by quantum error correction [8] to realize a fault-tolerant quantum computer (FTQC) [9].
@@
The surface codes [10, 11] are one of the most promising approaches for a fault-tolerant quantum computer using superconducting qubits, since they can be implemented on a two-dimensional square lattice and have high noise resilience [12, 13]. In FTQC whole computation has to be performed fault-tolerantly, with quantum information encoded in the logical qubits. While Clifford gates are relatively easy to do so, non-Clifford gates, such as the $T$ gate, are hard to execute fault-tolerantly [14]. Therefore, magic state distillation (MSD) [15] is employed to prepare a high fidelity magic state $T\ket{+}$ from noisy ones, which is hence used to implement the $T$ gate via gate teleportation [16].
@@
While MSD is a crucial operation for achieving universal quantum computation, it requires a large number of qubits, which is an obstacle to the realization of FTQC [17]. This is because in most MSD protocols distillation is performed by using logical qubits with concatenating QEC codes which have the transversality of the $H$ gate or $T$ gate. In order to mitigate this, a physical level distillation protocol has been proposed with error detection [18]. While this approach has great potential to reduce the physical overhead, it has not been used for resource estimation because of the lack of concrete implementation on an architecture compatible with the surface codes.
@@
In this work, we propose zero-level distillation to prepare a high fidelity logical magic state using physical qubits and nearest-neighbor two-qubit gates on a square lattice. At first, a magic state is encoded into the Steane code non-fault-tolerantly. Then such a noisy magic state is verified by using the Hadamard test of the logical $H$ gate. The logical magic state is teleported into the rotated or planar surface codes. In addition to the teleportation-based approach, we also develop a fault-tolerant code conversion from the Steane code to the surface code, which allows us to save the number of qubits further while the depth is increased.
@@
Throughout these processes, the operations employed are single-qubit gates/measurements and nearest-neighbor two-qubit gates. Furthermore, any single-point error is detected by the syndrome measurements and Hadamard test, and hence the logical error rate scales $\mathcal{O}(p^{2})$ in the physical error rate $p$. Since we do not need QEC codes for fault-tolerant logical Clifford gates, the spacial overhead is much smaller than the conventional MSDs. For a performance analysis, we fully simulate zero-level distillation circuits consisting of approximately 40 qubits reducing the number of qubits required for simulation to 23 qubits. As a result, zero-level distillation reduces the logical error rate of a magic state $p_{L}$ to $p_{L}\simeq 100\times p^{2}$ by only using 25 physical depths for the rotated surface code (23 for the planar surface code) For example, in the case of $p=10^{-4}$ ($p=10^{-3}$), the logical error rate results in $p_{L}=10^{-6}$ ($p_{L}=10^{-4}$), while the success probability is fairly high,  70% (95%). In the context of early-FTQC architecture [19], the zero-level distillation facilitates a significant enhancement, increasing the reliability of non-Clifford gate operations by an order of magnitude of two, in contrast to a scenario without MSD. In the context of full-fledged FTQC, zero-level distillation combined with conventional multi-level distillation [20] allows us to save the number of physical qubits significantly to achieve a given accuracy by reducing the number of levels.
@@
The rest of the paper is organized as follows. In Sec II, we provide a preliminary explanation of the Steane code and MSD. In Sec III, we provide a detailed description of our proposal, zero-level distillation. In Sec IV, we explain the protocol to convert the Steane code state to the rotated surface code directly. In Sec V, we present the outcomes of our numerical simulation. Then, Sec VI is devoted to a conclusion.
@@
In this section, we preliminarily explain the Steane code and MSD based on the transversality of the Hadamard gate on it.
The Steane code is a [[7,1,3]] stabilizer code, which can correct an arbitrary single-qubit Pauli error and detect an arbitrary two-qubit Pauli error. Stabilizer generators are defined by $XIXIXIX,XXIIXXI,XIXXIXI,\$$ZIZIZIZ,ZZIIZZI,ZIZZIZI$, each of them corresponds to three colored faces in Fig. 1. On the Steane code, all Clifford gates can be implemented transversally, while non-Clifford gates such as the $T$ gate cannot. However, by using the transversality of the Hadamard $H$ gate, a special resource state, the so-called magic state can be prepared fault-tolerantly as follows.
@@
A magic state |$\ket{A}$ is defined by |$\ket{A}\equiv e^{-i\frac{\pi}{8}Y}\ket{+}$, which is an eigenstate of $H$. By using the transversality of $H$ gate and controlled-Hadamard gates $\Lambda(H)$, we can implement a Hadamard test of the logical Hadamard operator followed by decoding as shown in Fig. 2 [21, 22], where the physical controlled-Hadamard gate $\Lambda(H)$ is constructed by $A^{\dagger}\Lambda(X)A$ from the CNOT gate $\Lambda(X)$ and the non-Clifford gate $A$. Suppose that all Clifford gates are ideal, and only non-Clifford gates are subject to errors with a probability $p$, then the above Hadamard test of the logical Hadamard operator and decoding process detects up to two errors and hence provides the magic state with error rate $O(p^{3})$. In order to realize ideal Clifford gates, each qubit in Fig. 2 is further encoded into a logical qubit, and error-corrected Clifford gates are used. However, this approach costs a large amount of physical qubits and operations due to the concatenation of two QEC codes.
A lower-cost distillation method with flag qubits has been proposed by Goto [18]. This protocol works with noisy Clifford gates by carefully designing the circuit using flag qubits so that any single point of error introduced by the Clifford gates never ruins the distillation. Since this approach does not require logical qubits for the Clifford gates, it has the potential to significantly reduce the number of qubits required for MSD. However, the original proposal in Ref. [18] is based on the Steane code and all-to-all gate connectivity, which makes it difficult to apply for FTQC with the surface code on a nearest-neighbor architecture.
@@
Figure 3: (Left) The zero-level distillation circuit. (Right) The detailed circuits for encoding of noisy magic state, preparation of cat state, and distillation are shown. The circuit in the blue box encodes a magic state encoded with the Steane code non-fault-tolerantly. The circuit in the red box prepares a cat state. The circuit in the green box is the distillation circuit utilizing the Hadamard test.
In order to reduce the overhead for magic state distillation, we propose zero-level distillation, which is a protocol to prepare a logical magic state on the surface code without using multiple logical qubits. We perform physical level distillation using the Steane code by carefully designing the distillation circuit fault-tolerantly with a smaller number of physical qubits and nearest-neighbor two-qubit gates on the square lattice. Then, the logical magic state is teleported from the Steane code to the surface code, which can be viewed as lattice surgery between color and surface codes [23]. (In the next section, we will also provide another approach based on a code conversion from the Steane code to the surface code without teleportation.) This combination of two QEC codes allows us to prepare the logical magic state using a smaller number of physical qubits with high fidelity on the square lattice as seen below in detail.
Zero-level distillation consists of three key processes, non-fault-tolerant magic state encoding, post-selection with the Hadamard test, and teleportation-based injection with lattice surgery. The detailed steps are as follows (see schematic circuit diagram in Fig. 3):
@@
    Encode a magic state in the Steane code non-fault-tolerantly. In parallel, a 7-qubit cat state is also prepared.
@@
    Measure the $Z$-stabilizers on the Steane code and the surface code.
@@
    Measure qubits on the Steane code directly in $X$-basis to complete teleportation.
@@
Fig. 3 shows the circuit for encoding noisy logical magic state and preparation of the cat state corresponding to step (i). The blue box in Fig. 3 shows the circuit for encoding a magic state non-fault-tolerantly with the Steane code. As shown in Fig. 4, the circuit can be constructed by using nearest-neighbor two-qubit gates on the square lattice. The red box in Fig. 3 shows the circuit for encoding a cat-state. While in the existing work [18] magic state distillation is executed using two ancilla qubits, the 7-qubit cat state is suitable in our situation where the qubit connectivity is very limited. Therefore, we use a 7-qubit cat state $\frac{1}{\sqrt{2}}(\ket{0}^{\otimes 7}+\ket{1}^{\otimes 7})$ as ancilla qubits. This allows data and ancilla qubits to be adjacent on a square lattice as shown in Fig. 4 and  5.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Figure 4: Qubits arrangement during encoding a magic state. Each dot represents a qubit. The blue dots and numbers correspond to the locations of the qubits shown in Fig. 3. The light blue arrows indicate the transfer of qubits using one-bit teleportations. The blue lines indicate the application of CNOT gates.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 5: Qubits arrangement during encoding a cat state. Each dot represents a qubit. The orange dots and numbers correspond to the locations of the qubits in Fig. 3. The yellow arrows indicate the transfer of qubits using one-bitteleportations. The red lines indicate the application of CNOT gates.
@@
In step (ii), the Hadamrd test for the logical Hadamard gate is performed by using the cat state. The green box in Fig. 3 shows the distillation circuit with the Hadamard test, where the controlled-Hadamard gate $\Lambda(H)$ is implemented as $A^{\dagger}\Lambda(X)A$. Fig. 6 (right) shows the details of the arrangement of qubits during distillation. Since data and ancilla qubits are located side by side, CNOT gates can be applied between them directly. After the logical controlled-Hadamard gate, the ancilla qubits from 7 to 13 in Fig. 3 are measured in the $X$ basis. If the parity of the measurement outcomes is odd indicating any error, the distillation process is rejected. Otherwise, the output is accepted and proceeds to the next step. Then some qubits are moved for the lattice surgery as shown in Fig. 7(LABEL:sub@fig:rotated_2).
@@
Figure 6: Qubits arrangement during the Hadamard test. The blue and orange dots and numbers indicate the locations of the qubits shown in Fig. 3. A magic state is encoded in the blue qubits and a cat state is encoded in the orange qubits. The green lines indicate the CNOT gates.
@@
In parallel, the |$\ket{+}_{L}$ state is encoded with the rotated surface code by preparing $\ket{+}$ states and measuring the $Z$-stabilizers as shown in Fig. 7(LABEL:sub@fig:rotated_1).
@@
In step (iii), Steane and rotated surface codes are merged as shown in Fig. 7(LABEL:sub@fig:rotated_3) and then split via the lattice surgery. This results in projection with respect to $L_{Z}^{\mathrm{Steane}}\otimes L_{Z}^{\mathrm{Surface}}$ since its eigenvalue can be obtained from the product of $Z$-type operators at the boundary shown by the purple area of Fig. 7(LABEL:sub@fig:rotated_3). During the lattice surgery, it is necessary to measure the $Z$-stabilizers twice to detect measurement errors. If the measurement outcomes differ, such an event is rejected, and the protocol must be restarted from the beginning.
@@
In step (iv), $Z$-stabilizers on the Steane code and the rotated surface code are measured. Fig. 8 illustrates the syndrome measurement circuits for the Steane code [24], where the ancilla qubits in Fig. 8 are represented by red, blue, or green dots in Fig. 7(LABEL:sub@fig:rotated_4). The measurements detect two-qubit errors introduced during merging, as well as errors during magic state encoding or |$\ket{+}_{L}$ state encoding.
@@
In step (v), each qubit in the Steane code is measured directly in the $X$ basis to obtain the eigenvalue of the $X$-stabilizers and logical operators as shown in Fig. 7(LABEL:sub@fig:rotated_4). Note that these eigenvalues should be interpreted appropriately according to the measurement outcome of the $X$-stabilizer at the boundary for splitting as usually done in the lattice surgery. This completes the teleportation of the distilled magic state on the Steane code to the rotated surface code. As usual, the Pauli frame of the logical qubit is updated according to the measurement outcome. Throughout all operations from the beginning, 25 physical depths are used. While we have demonstrated the proposed protocol for the rotated surface code, the planar surface code version is also shown in Appendix A.
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Figure 7: Qubits arrangement and measurements during teleportation from the Steane code to the rotated surface code. The white dots and numbers indicate the locations of the qubits shown in Fig. 3, where the distilled magic state is encoded in the white qubits. The light blue arrows indicate the transfer of qubits using teleportation. The red, blue and green faces indicate the stabilizers of the Steane code, and the orange and yellow faces indicate the stabilizers of the rotated surface code. The purple faces indicate the $Z$-stabilizers for lattice surgery. The red, blue, and green dots are the ancilla qubits for the stabilizer measurements. The gray lines indicate logical operators of the Steane code. The purple box and the gray box indicate logical operators of the rotated surface code.
Refer to caption Refer to caption
@@
Figure 8: The syndrome measurement circuits for the Steane code use three ancilla qubits (left) and two ancilla qubits (right), which correspond to red and blue/green in Fig. 7(LABEL:sub@fig:rotated_4).
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 9: Qubits arrangement and measurements during conversion from the Steane code to the rotated surface code. The blue and orange dots and numbers correspond to the locations of the qubits shown in Fig. 3. The blue arrows indicate the transfer of qubits using one-bit teleportations. The red, blue,and green faces indicate the stabilizers of the Steane code, and the orange and yellow faces indicate the stabilizers of the rotated surface code. The red, blue, and green dots are the ancilla qubits for the stabilizer measurements of the Steane code.

### v2

# Efficient Magic State Distillation by Zero-Level Distillation

@@ Abstract
Magic state distillation (MSD) is an essential element for universal fault-tolerant quantum computing, which distills a high-fidelity magic state from noisy magic states using ideal (error-corrected) Clifford operations. For ideal Clifford operations, it needs to be performed on the logical qubits and hence incurs a large spatiotemporal overhead, which is one of the major bottlenecks for the realization of fault-tolerant quantum computers (FTQCs). Here we propose zero-level distillation, which prepares a high-fidelity logical magic state at the physical level, namely zero level, using physical qubits and nearest-neighbor two-qubit gates on a square lattice. We develop a zero-level distillation circuit and show that distillation can be made even more efficient than the conventional sophisticated approaches with logical level distillations. The key idea involves the Knill et al.-type distillation using the Steane code and its careful mapping to the square-lattice architecture with error detection. The distilled magic state on the Steane-code state is then teleported or converted to surface codes. We numerically find that the error rate of the logical magic state scales as approximately $100\times p^{2}$ in terms of the physical error rate $p$. For example, with a physical error rate of $p=10^{-4}$ ($10^{-3}$), the logical error rate is reduced to $p_{L}=10^{-6}$ ($10^{-4}$), resulting in an improvement of 2 (1) orders of magnitude. This contributes to reducing both space and time overhead for early FTQC as well as full-fledged FTQC combined with conventional multilevel distillation protocols.
@@ I Introduction
Quantum computers are expected to provide advantages in solving problems that are intractable for classical computers such as prime factorization [1], linear system solver [2], and quantum chemistry [3]. Significant experimental efforts have been dedicated to the realization of quantum computers based on various physical systems. Among these, the superconducting system is one of the most promising candidates, and systems with 50-100 qubits have already been experimentally demonstrated [4, 5]. These quantum computers are called noisy intermediate-scale quantum computers (NISQ) [6], whose noise level is still high and the number of qubits is still limited, are currently unable to run sophisticated quantum algorithms with theoretically proven quantum speedup. Although NISQ-aware algorithms are being developed [7], an ultimate solution to these problems is to protect quantum information through quantum error correction [8] to realize a fault-tolerant quantum computer (FTQC) [9].
@@
Surface codes [10, 11] are one of the most promising approaches for a fault-tolerant quantum computer using superconducting qubits, since they can be implemented on a two-dimensional square lattice and have high noise resilience [12, 13]. In FTQC, the entire computation has to be performed fault tolerantly, with quantum information encoded in logical qubits. While Clifford gates can be implemented relatively easily in a fault-tolerant manner, non-Clifford gates, such as the $T$ gate, are hard to execute fault tolerantly [14]. Therefore, magic state distillation (MSD) [15] is employed to prepare a high-fidelity magic state $T\ket{+}$ from noisy ones, which is hence used to implement the $T$ gate via gate teleportation [16].
@@
While MSD is a crucial operation for achieving universal quantum computation, it requires a large number of qubits, which is an obstacle to the realization of FTQC [17]. This is because in most MSD protocols distillation is performed by using logical qubits with concatenating a QEC code, which features transversal implementations of the $H$ or $T$ gates. In order to mitigate this, physical-level distillation protocols have been proposed with error detection [18, 19, 20]. Physical-level distillation has great potential to reduce the physical overhead, and its effectiveness has been demonstrated in an experiment with trapped ions [21]. However, this physical-level approach has not been used for resource estimation because it employs all-to-all or complicated gate connectivity, and a concrete implementation on an architecture compatible with surface codes is lacking.
@@
In this work, we propose zero-level distillation to prepare a high-fidelity logical magic state using physical qubits and nearest-neighbor two-qubit gates on a square lattice. Here zero level refers to performing distillation entirely at the physical level instead of distillation at a higher level using logical qubits. More precisely, in zero-level distillation, a magic state is initially encoded into the Steane code non-fault-tolerantly. Then this noisy magic state is verified using the Hadamard test of the logical $H$ gate following the Knill et al.-type protocol [22]. The logical magic state is teleported to the rotated or planar surface code. In addition to the teleportation-based approach, we also develop a fault-tolerant code conversion from the Steane code to the surface code, which allows us to reduce the number of qubits further while the depth is increased.
Throughout these processes, the operations employed are single-qubit gates and measurements and nearest-neighbor two-qubit gates. Furthermore, any single-point error is detected by syndrome measurements and the Hadamard test, and hence the logical error rate scales as $\mathcal{O}(p^{2})$ with respect to the physical error rate $p$. Since QEC codes are not needed for fault-tolerant logical Clifford gates, the spatial overhead is much smaller than that of conventional MSD protocols.
@@
It might be thought that the advantage of the proposed zero-level distillation is limited to be the case with very low physical error rate because of the square-lattice constraint. However, this is not the case if the distillation circuit is carefully designed. For performance analysis, we fully simulate zero-level distillation circuits consisting of approximately 40 qubits, reducing the number of qubits required for simulation to 23 qubits. As a result, zero-level distillation reduces the logical error rate of a magic state $p_{L}$ to $p_{L}\simeq 100\times p^{2}$ by using only a physical depth of 25 for the rotated surface code (23 for the planar surface code). For example, in the case of $p=10^{-4}$ ($p=10^{-3}$), the logical error rate results in $p_{L}=10^{-6}$ ($p_{L}=10^{-4}$), while the success rate is fairly high, 70% (95%). Although the $100p^{2}$ scaling is slightly worse than $35p^{3}$ for the conventional method, it offers practical advantages in terms of the overall spatiotemporal overhead. Specifically, in the context of the early-FTQC architecture [23], the zero-level distillation facilitates a significant enhancement, increasing the reliability of non-Clifford gate operations by 2 orders of magnitude, in contrast to a scenario without MSD. For full-fledged FTQC, zero-level distillation combined with conventional multilevel distillation [24] allows us to save the number of physical qubits significantly to achieve a given accuracy by reducing the number of levels.
@@
The rest of the paper is organized as follows. In Sec. II, we provide a preliminary explanation of the Steane code and MSD. In Sec. III, we provide a detailed description of our proposal, zero-level distillation. In Sec. IV, we explain the protocol to convert the Steane-code state to the rotated surface code directly. In Sec. V, we present the outcomes of our numerical simulation. In Sec. VI, we discuss the implications of zero-level distillation to early-FTQC and full-fledged FTQC. Then, Sec. VII is devoted to a conclusion.
@@ II PRELIMINARY
In this section, we provide a preliminary explanation of the Steane code and MSD based on the transversality of the Hadamard gate on it.
The Steane code is a $\llbracket 7,1,3\rrbracket$ stabilizer code, which can correct an arbitrary single-qubit Pauli error or detect an arbitrary two-qubit Pauli error. Stabilizer generators are defined by $XIXIXIX,XXIIXXI,XIXXIXI,\$$ZIZIZIZ,ZZIIZZI,ZIZZIZI$, each corresponds to one of the three colored faces in Fig. 1. On the Steane code, all Clifford gates can be implemented transversally, while non-Clifford gates such as the $T$ gate cannot. Nonetheless, leveraging the transversality of the Hadamard $H$ gate, a special resource state, referred to as the magic state, can be prepared fault tolerantly as follows.
@@
A magic state |$\ket{A}$ is defined as |$\ket{A}\equiv e^{-i\frac{\pi}{8}Y}\ket{+}$, which is an eigenstate of $H$. Leveraging the transversality of $H$ gate and controlled-Hadamard gates $\Lambda(H)$, we implement a Hadamard test of the logical Hadamard operator followed by decoding, as shown in Fig. 2 [22, 25], where the physical controlled-Hadamard gate $\Lambda(H)$ is constructed by $A^{\dagger}\Lambda(X)A$ using the CNOT gate $\Lambda(X)$ and the non-Clifford gate $A$. Suppose that all Clifford gates are ideal, and only non-Clifford gates are subject to errors with a probability $p$, the above Hadamard test of the logical Hadamard operator and decoding process detect up to two errors. This provides the magic state with error rate $O(p^{3})$. In order to make Clifford gates ideal, each qubit in Fig. 2 is further encoded into a logical qubit, and error-corrected Clifford gates are employed for distillation. However, this approach costs a large amount of physical qubits and operations due to the concatenation of two QEC codes.
A lower-cost distillation method with flag qubits has been proposed by Goto [18]. This protocol works with noisy Clifford gates by carefully designing the circuit with flag qubits so that any single-point error during the Clifford gates does not compromise the entire distillation process. Since this approach eliminates the need for logical qubits in Clifford gates, it can substantially reduce the number of qubits needed for MSD. However, the original proposal in Ref. [18] relies on the Steane code and all-to-all gate connectivity, which makes it difficult to apply for FTQC with the surface code on a nearest-neighbor architecture.
@@ III ZERO-LEVEL DISTILLATION
Figure 3: (a) The zero-level distillation circuit. (b) The detailed circuits for encoding of noisy magic state, preparation of cat state, and distillation. The circuit in the blue box encodes a magic state encoded with the Steane code non-fault-tolerantly. The circuit in the red box prepares a cat state. The circuit in the green box is the distillation circuit utilizing the Hadamard test.
In order to reduce the overhead of magic state distillation, we propose zero-level distillation. This protocol prepares a logical magic state on the surface code without using multiple logical qubits. We achieve physical-level distillation using the Steane code by carefully designing a fault-tolerant distillation circuit with fewer physical qubits and nearest-neighbor two-qubit gates on a square lattice. Then, the logical magic state is teleported from the Steane code to the surface code, which can be viewed as lattice surgery between color and surface codes [26]. (In the next section, we will also provide an alternative approach based on a code conversion from the Steane code to the surface code without teleportation.) This combination of two QEC codes allows us to prepare the logical magic state using fewer physical qubits with high fidelity on a square lattice as detailed below.
Zero-level distillation consists of three key processes: non-fault-tolerant magic state encoding, postselection via the Hadamard test, and teleportation-based injection using lattice surgery. The detailed steps are as follows (see the schematic circuit diagram in Fig. 3):
@@
    Encode a magic state in the Steane code non-fault-tolerantly. In parallel, a seven-qubit cat state is also prepared.
@@
    Measure the $Z$ stabilizers on the Steane code and the surface code.
@@
    Measure qubits on the Steane code directly in $X$ basis to complete teleportation.
@@
Fig. 3 shows the circuit for encoding a noisy logical magic state and preparing the cat state corresponding to step (i). The blue box in Fig. 3 shows the circuit for non-fault-tolerant encoding of a magic state using the Steane code. As shown in Fig. 4, the circuit can be constructed by using nearest-neighbor two-qubit gates on a square lattice.
The red box in Fig. 3 shows the circuit for preparing a cat state. While in previous work [18], magic state distillation is performed using two ancilla qubits, a seven-qubit cat state is more suitable in our situation where the qubit connectivity is highly limited. Therefore, we use a seven-qubit cat state $\frac{1}{\sqrt{2}}(\ket{0}^{\otimes 7}+\ket{1}^{\otimes 7})$ as ancilla qubits. This allows data and ancilla qubits to be adjacent on a square lattice as shown in Figs. 4 and  5.
Note that in Figs. 4 and  5, some qubits have to be moved so that CNOT gates between neighboring qubits can be performed. Instead of using the swap operation, we employed one-bit teleportation to move a qubit to the neighboring site. This is simply because the swap operation requires three CNOT gates, which results in a greater circuit depth. However, if a swap gate can be implemented quickly and reliably, such as the $i$SWAP gate, it can be replaced with such a swap operation.
@@
Figure 4: Qubit arrangement and stabilizer deformation during encoding a magic state. The left figure represents the $X$-stabilizer deformation, and the right figure illustrates the $Z$-stabilizer deformation. Each dot represents a qubit: the blue dots and numbers correspond to the locations of the qubits shown in Fig. 3, and the light blue arrows indicate the transfer of qubits using one-bit teleportations. The blue bold edges indicate the CNOT gates, where the $+$ symbol on one side represents the target qubit.
@@
Figure 5: Qubit arrangement during encoding a cat state. Each dot represents a qubit: the orange dots and numbers correspond to the locations of the qubits in Fig. 3. The yellow arrows indicate the transfer of qubits using one-bit teleportations. The red bold edges indicate the CNOT gates where the $+$ symbol on one side represents the target qubit.
@@
In step (ii), a Hadamard test for the logical Hadamard gate is performed using the cat state. The green box in Fig. 3 shows the distillation circuit with the Hadamard test, where the controlled-Hadamard gate $\Lambda(H)$ is implemented as $A^{\dagger}\Lambda(X)A$. Fig. 6 (right) shows the details of the arrangement of qubits during distillation. Since the data and ancilla qubits are positioned adjacent to each other, CNOT gates can be applied directly between them. After the logical controlled-Hadamard gate, the ancilla qubits from 7 to 13 in Fig. 3 are measured in the $X$ basis. If the parity of the measurement outcomes is odd indicating any error, the distillation process is rejected. Otherwise, the output is accepted and proceeds to the next step. Next, some qubits are repositioned for the lattice surgery as shown in Fig. 7(b).
@@
Figure 6: Qubit arrangement during the Hadamard test. The blue and orange dots and numbers indicate the locations of the qubits shown in Fig. 3. A magic state and a cat state are encoded in the blue and orange qubits, respectively. The green edges indicate the CNOT gates, where the $+$ symbol on one side represents the target qubit.
@@
Simultaneously, the |$\ket{+}_{L}$ state is encoded using the rotated surface code by preparing $\ket{+}$ states and measuring the $Z$ stabilizers as shown in Fig. 7(a).
@@
In step (iii), Steane and rotated surface codes are merged as shown in Fig. 7(c), and then split via the lattice surgery. This results in a projection onto $L_{Z}^{\mathrm{Steane}}\otimes L_{Z}^{\mathrm{Surface}}$ as its eigenvalue can be obtained from the product of $Z$-type operators at the boundary, indicated by the purple area of Fig. 7(c). During the lattice surgery, the $Z$ stabilizers must be measured twice to detect measurement errors. If the measurement outcomes differ, the event is rejected, and the protocol must be restarted from the beginning.
@@
In step (iv), $Z$ stabilizers on the Steane code and the rotated surface code are measured. Fig. 8 illustrates the syndrome measurement circuits for the Steane code [27], where the ancilla qubits in Fig. 8 correspond to red, blue, or green dots in Fig. 7(d). The measurements detect two-qubit errors introduced during merging, as well as errors during magic state encoding or |$\ket{+}_{L}$-state encoding.
@@
In step (v), each qubit of the Steane code is directly measured in the $X$ basis to obtain the eigenvalues of the $X$ stabilizers and logical operators as shown in Fig. 7(d). Note that these eigenvalues should be interpreted appropriately according to the measurement outcome of the $X$ stabilizer at the boundary for splitting as usually done in the lattice surgery. This completes the teleportation of the distilled magic state from the Steane code to the rotated surface code. As usual, the Pauli frame of the logical qubit is updated based on the measurement outcome. Throughout the entire procedure, a physical depth of 25 is used. While we have demonstrated the proposed protocol using the rotated surface code, an implementation with the planar surface code is also provided in Appendix A.
@@
Figure 7: Qubit arrangement and measurements during teleportation from the Steane code to the rotated surface code. The white dots and numbers indicate the locations of the qubits shown in Fig. 3, where the distilled magic state is encoded in the white qubits. The light blue arrows indicate the transfer of qubits using teleportation. The red, blue and green faces indicate the stabilizers of the Steane code, and the orange and yellow faces indicate the stabilizers of the rotated surface code. The purple faces indicate the $Z$ stabilizers for lattice surgery. The red, blue, and green dots are the ancilla qubits for the stabilizer measurements. The gray lines indicate logical operators of the Steane code. The purple box and the gray box indicate logical operators of the rotated surface code.
@@
Figure 8: (a) The syndrome measurement circuit for the Steane code using three ancilla qubits. (b) The syndrome measurement circuit using two ancilla qubits. They correspond to red and blue-green in Fig. 7(d).
The above zero-level distillation generates a magic state encoded in a $d=3$ surface code, but it is desirable that the output state is kept on the surface code with a larger code distance to avoid error accumulation. Therefore, the surface code is further expanded fault tolerantly as shown in Fig. 9. In this process, |$\ket{0}$ and $\ket{+}$ ancilla qubits are properly prepared [23] and stabilizer measurements are performed.
Figure 9: An expansion of the code distance from $d=3$ to $d=5$ in the rotated surface code. (a) Before expansion. (b) The output magic state on the $d=3$ surface code is located in the corner and other data qubits are initialized to |$\ket{0}$ or $\ket{+}$ ancilla qubits. Then syndrome measurements are done, where error correction is performed without postselection.
@@
Figure 10: Qubit arrangement and measurements during conversion from the Steane code to the rotated surface code. The blue and orange dots and numbers correspond to the locations of the qubits shown in Fig. 3. The blue arrows indicate the transfer of qubits using one-bit teleportations. The red, blue, and green faces indicate the stabilizers of the Steane code, and the orange and yellow faces indicate the stabilizers of the rotated surface code. The red, blue, and green dots are the ancilla qubits for the stabilizer measurements of the Steane code.